rods»view/3

Status Summary for ROD ID 3

<<Prev | Next>>

Test NameStatusNotesModified
Board Status 0 showed many broken channels between V2 and top V4; board taken to Israel where new V2 was mounted problems between V2 and V4 are fixed but now evidence of (one, two) broken channels between VME chip and V4 top (pins 33 and 2 - was just pin 33 in Israel, back at CERN appears 2 has a problem as well) pushing on top V4 HARD can fix this problem - clears both 2 channels that show bad connection (13 June 2007) board fixed by CERN shop; however, 101 on test fifo data lines 2(0)-3(31) shows 111 instead of 101 - need to mask off bottom two GOLs to get data through from test fifo; normal data works fine June 28: neither pushing on the test fifo nor on the V2 fixes the problem 4 July: pusing on the test fifo quite hard (and in the right place) fixes the problem 5 July: board brought to CERN shop for reheating of test fifo 9 July: some soft short? Consumes too much power on bench on 3.3 V supply and does not work with VME crate 12 July: with Rick's help, seen that power draw results from V2; pushing on V2 increases power consumption; should replace or remove or reball V2?? 15 August 2007: conn_test -> 2 chnls bad betw. V2/V4b; 2 chnls bad betw. V2/V4t; 9 chnls stuck HIGH betw. V2/V4t; 2 chnls bad betw. V4t/V (VME_chip) Consumes too small power on bench. Sent to AppliCAD for temperature profiling. 24-06-2011
fuses 1
12-12-2007
U31
12-12-2007
U32
12-12-2007
U33
12-12-2007
U34
12-12-2007
U35
12-12-2007
U81
12-12-2007
U82
12-12-2007
I5V_initial
12-12-2007
I3V initial
12-12-2007
proms load
12-12-2007
I5V proms
12-12-2007
I3V proms
12-12-2007
V4T temp
12-12-2007
V4B temp
12-12-2007
connectivity
12-12-2007
loop data
12-12-2007
slink
12-12-2007
dram
12-12-2007
nse communication
12-12-2007
sram
12-12-2007
optical loop
12-12-2007
patch panel
12-12-2007
qpll lock
12-12-2007
ttc connectivity
12-12-2007
test fifo 0
See history
12-12-2007
ddr reset
12-12-2007
assembly 0
see history.
12-12-2007